OpenStudy (anonymous):

Verify the identity. sin (x + y) - sin (x - y) = 2 cos x sin y

hartnn (hartnn):

do you know the formula for sin(A+B) =... ?

OpenStudy (anonymous):

Sin(a+b) = sin a . cos b + Cos a. Sin b sin (a-b) = sin a . cos b - Cos a. Sin b

Parth (parthkohli):

Yeah, that's pretty much it ^